The double of Mr. beery's age squared is equal to his age times 67 plus 105. How old is Mr. Beery?

Answer by ankor@dixie-net.com(22740)

Let a = his age
write an equation for exactly what it says
:
The double of Mr. beery's age squared is equal to his age times 67 plus 105.
2a^2 = 67a + 105
A quadratic equation
2a^2 - 67a - 105 = 0
You can use the quadratic formula but this will factor to
(2a+3)(a-35) = 0
The positive solution
a = 35 is Mr B's age